# Attacks on fairness in Federated Learning

### v2

@@ 1 Introduction
In this paper we introduce the concept of attacks on fairness in Federated Learning. A similar idea has previously been investigated in the context of centralised models [8, 9, 10]. However, the threat model in FL is significantly different because we are constrained by controlling a relatively small portion of the clients used to train the FL model. In order for FL systems to become resilient to fairness attacks, a very different set of defences will need to be implemented compared to in the centralised case. We discuss these differences further in Section 4.2.
@@
  • We introduce the concept of attacking fairness in FL and explain the threat model (Section 2.2).
@@
  • We discuss how existing FL backdoor defences could be modified to defend models against fairness attacks (Section 4.3).
@@ 2.1 Fairness in Federated Learning
Previous work has discussed at length the concept of fairness in FL. [11] describes three different types of fairness in FL systems. In this paper we are interested in compromising attribute level fairness. That is, our attack aims to influence the model to create a significant discrepancy between the expected accuracy of data points with the targeted attribute and the overall test set accuracy. We discuss further the relationship of our work, to that aimed at the centralized setup in Section 4.2.
@@ 2.3 Attacks on Fairness
where we control both $n_{0}$ and $\mathbf{v}$. In practice, we can only estimate the values of $\mathbf{u}_{i}$ by predicting updates with our own representative dataset. Figure 1 shows a visual representation of how the FedAvg aggregation will reproduce our target update $\mathbf{m}$.
@@ 3 Evaluation
Table 1 presents the results from our testing. We tested two configurations of our attack. The bottom row has the attack performed from the very first training round, while the middle row trained the model on only clean clients for the first 80 rounds. The table shows that our attack has been successful at producing a difference in accuracy between the two targeted classes and the others. Both attack configurations were successful, showing that our attack would work on an existing system, where it was not introduced from the first training round.
@@
Figure 2 shows the accuracy of each class per round on the UCI census dataset. We can see that after the attack began on round 10, the target class accuracies immediately becomes higher than that of other classes, as expected.

### v3

@@ 1 Introduction
A substantially updated version of this paper can be found under the title “Can Private Machine Learning Be Fair?". Link
This version of the paper has been extensively summarized to fit the page limit of NeurIPS camera ready, and some materials, experiments, discussions, and methods are moved to appendix, which might make some parts hard to follow or cause inconsistencies. To avoid such cases, please read our arXiv version instead [1]
@@
In this paper we introduce the concept of attacks on fairness in Federated Learning. A similar idea has previously been investigated in the context of centralised models [8, 9, 10]. However, the threat model in FL is significantly different because we are constrained by controlling a relatively small portion of the clients used to train the FL model. In order for FL systems to become resilient to fairness attacks, a very different set of defences will need to be implemented compared to in the centralised case. We discuss these differences further in Section˜4.2.
@@
  • We introduce the concept of attacking fairness in FL and explain the threat model (Section˜2.2).
@@
  • We discuss how existing FL backdoor defences could be modified to defend models against fairness attacks (Section˜4.3).
@@ 2.1 Fairness in Federated Learning
Previous work has discussed at length the concept of fairness in FL. [11] describes three different types of fairness in FL systems. In this paper we are interested in compromising attribute level fairness. That is, our attack aims to influence the model to create a significant discrepancy between the expected accuracy of data points with the targeted attribute and the overall test set accuracy. We discuss further the relationship of our work, to that aimed at the centralized setup in Section˜4.2.
@@ 2.3 Attacks on Fairness
where we control both $n_{0}$ and $\mathbf{v}$. In practice, we can only estimate the values of $\mathbf{u}_{i}$ by predicting updates with our own representative dataset. Figure˜1 shows a visual representation of how the FedAvg aggregation will reproduce our target update $\mathbf{m}$.
@@ 3 Evaluation
Table˜1 presents the results from our testing. We tested two configurations of our attack. The bottom row has the attack performed from the very first training round, while the middle row trained the model on only clean clients for the first 80 rounds. The table shows that our attack has been successful at producing a difference in accuracy between the two targeted classes and the others. Both attack configurations were successful, showing that our attack would work on an existing system, where it was not introduced from the first training round.
@@
Figure˜2 shows the accuracy of each class per round on the UCI census dataset. We can see that after the attack began on round 10, the target class accuracies immediately becomes higher than that of other classes, as expected.